OpenStudy (coexist2):

Quilt squares are cut on the diagonal to form triangular quilt pieces. The hypotenuse of the resulting triangles is 8 inches long. What is the side length of each piece?

OpenStudy (geliraymundo):

The diagonal of a square is equal to \[side \sqrt{2}\] So, \[8=side \sqrt{2}\] so side would be \[\frac{ 8 }{\sqrt{2} }\] simplifying it by multiplying both the numerator and denominator with \[\sqrt{2}\] Making it \[4\sqrt{2}\]

OpenStudy (coexist2):

thank you for the help.